# Discovering Global False Negatives On the Fly for Self-supervised Contrastive Learning

### v1

@@ 4.2 Bimodal Experiments
Experiment Setup. Following previous work (Wei et al., 2024), we use a 12-layer transformer (Vaswani et al., 2017) as the text encoder, and ResNet50 as the vision encoder. All experiments are conducted in a multi-node setting with 2 nodes, each with two A100 40GB GPUs. We pretrain for 37 epochs with a global batch size of 1024. For SogCLR, we start using FNC/GloFND after 15 epochs, setting $\alpha=5e-4$, while for FastCLIP we start after 20 epochs with $\alpha=1e-3$. For both losses, we initialize $\lambda_{i}=1$ and use Adam updates with a learning rate of 0.05. More details on hyperparameters can be found in Appendix H.
@@ 4.3 Ablation Study
Impact of Starting Epoch. GloFND requires a sufficiently pre-trained network to ensure that the similarity between embeddings reflects semantic similarity. This is achieved by applying GloFND after a certain number of training epochs. However, the optimal start time involves a trade-off. If GloFND is applied too early, the embedding space may not be well-formed, leading to incorrect classification of pairs as false negatives. Conversely, if applied too late, the potential benefits of GloFND may be limited due to insufficient training time. To assess the necessity and impact of this “wait” period, we evaluate GloFND’s semi-supervised performance on ImageNet100 as we vary the start epoch in $\{10,30,50,70,90,110\}$. The linear evaluation results, presented in LABEL:fig:warmup_100, show that GloFND’s performance improves as the wait time increases, peaking at 70 epochs. Beyond this point, performance begins to decline.
@@ 5 Conclusions
Limitations. Since the focus of this paper is on false negative detection for contrastive learning, we address false negatives through filtering. While this straightforward approach has proven effective in our settings, future work could explore more advanced methods that may further enhance downstream performance. Additionally, the benefits of GloFND and similar false-negative techniques on downstream tasks depend on how false negatives are defined within the downstream task and the proportion of false negatives in the pretraining dataset, a factor we expect to become more significant as foundational model datasets continue to scale.
B More Details on Extension to Bimodal CL
@@
Our approach GloFND can be extended to resolve the global false negative discovery in bimodal CL, e.g., CLIP [Radford et al., 2021]. Consider a dataset of image-text pairs $\mathcal{D}=\{(\mathbf{x}_{1},\mathbf{t}_{1}),\dotsc,(\mathbf{x}_{n},\mathbf{t}_{n})\}$, a collection of image augmentation operators $\mathcal{P}_{I}$, and a collection of text augmentation operators $\mathcal{P}_{T}$. Suppose that the encoder network $E_{I}$ for images and the encoder network $E_{T}$ for text are parametrized by $\mathbf{w}$. For each $(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}$, the negative dataset for each anchor image $\mathbf{x}_{i}$ is $\mathcal{S}_{I,i}^{-}=\{A_{T}(\mathbf{t})\mid\forall A_{T}\in\mathcal{P}_{T},\forall(\mathbf{x},\mathbf{t})\in\mathcal{D}\}$ while the negative dataset for each anchor text $\mathbf{t}_{i}$ is $\mathcal{S}_{T,i}^{-}=\{A_{I}(\mathbf{x})\mid\forall A_{I}\in\mathcal{P}_{I},\forall(\mathbf{x},\mathbf{t})\in\mathcal{D}\}$. For the $i$-th anchor image $\mathbf{x}_{i}$ with representation $\mathbf{z}_{I,i}$, the threshold $\lambda_{I,i}\in[-1,1]$ for finding the top $\alpha$text neighbors among all negatives $\mathcal{S}_{I,i}^{-}$ can be solved by
$\min_{\nu\in[-1,1]}\nu\alpha+\frac{1}{|\mathcal{S}_{I,i}^{-}|}\sum_{\mathbf{t}\in\mathcal{S}_{I,i}^{-}}\left(\mathrm{sim}(\mathbf{z}_{I,i},E_{T}(\mathbf{t}))-\nu\right)_{+}.$
@@
Similarly, we can obtain the threshold $\lambda_{T,i}\ in[-1,1]$ for the $i$-th anchor text $\mathbf{t}_{i}$. Given the thresholds $\bm{\lambda}_{I},\bm{\lambda}_{T}$, the bimodal contrastive loss can be written as
@@
C More Details on GloFND
Appendix D High-level Intuition for $\alpha$ Hyperparameter
@@
The hyperparameter $\alpha$ allows GloFND to adapt to different definitions of false negatives. Note that the ”optimal definition” of false negative depends on the task of interest. For instance, in a dataset like ImageNet, we might consider two classification tasks: (1) classifying between cars and animals, and (2) classifying dog breeds. Two images of a dog from different breeds might be considered a false negatives for (1) but not for (2). Consequently, the optimal percentage of false negatives (i.e., $\alpha$) for tasks (1) and (2) is different. Thus, $\alpha$ enables GloFND to adapt to varying levels of granularity in false negative definitions.
@@
The value of $\alpha$ can be set based on prior knowledge of the number/probability of false negatives, the desired granularity of representations, or tuned like other hyperparameters such as the temperature. If $\alpha$ is too low, GloFND may fail to identify sufficient false negatives, leading to minimal impact on the learned representations, though not degrading performance, as setting $\alpha=0$ is equivalent to disabling GloFND. Conversely, if $\alpha$ is too high, GloFND may identify too many false negatives. If these are filtered out during training, the reduced number of negative pairs can limit contrastive learning, potentially harming performance. For tuning, it is recommended to start with a low $\alpha$ and gradually increase it until no further performance gains are observed.
Appendix E More Details on Computational Efficiency
When using GloFND with a contrastive method based on embedding similarity (e.g., SimCLR, SogCLR, and CLIP), pairwise similarities are already computed as part of the loss function. Thus, the only additional computations required are: (1) updating the $\lambda_{i}$ values for the mini-batch samples and (2) filtering the false negatives. Both operations involve basic matrix computations and run in linear time with respect to the number of pairs in a batch ($O(B^{2})$, where $B$ is the batch size). The computational overhead of GloFND is minimal compared to the cost of cosine similarity computation and forward/backpropagation.
Our experiments on ImageNet100 with a batch size of 128 show that GloFND introduces only a 2% increase in per-epoch training time for SogCLR (435.19 s for SogCLR + GloFND vs. 426.67 s for SogCLR). This overhead is comparable to that of the batch-wise method FNC (434.06 s).
F More Details on Experiments
@@
All the experiments are implemented using the PyTorch [Paszke et al., 2019] library. The unimodal experiments are run on a single NVIDIA A30 with 24GB memory size, while the bimodal experiments make use of a multi-node setup with 2 nodes, each with 2 NVIDIA A100 GPUs with 40GB each. The estimated amount of time to run a single experiment is 1.5 days for ImageNet100, and 14 hours for CC3M.
Appendix G Additional Details for Unimodal Experiments
Experiment Setup.
@@ Appendix G Additional Details for Unimodal Experiments
Following prior work [Yuan et al., 2022], we pretrain a ResNet-50 [He et al., 2015] with a 2-layer $128\times 128$ projection head on top of the backbone encoder. We use square root learning rate scaling ($0.075\times\sqrt{\text{BatchSize}}$) with a cosine decay schedule without restart. Additionally, we apply a linear learning rate warm-up for 10 epochs, where the learning rate linearly increases to its maximum value.
@@
We adopt the same augmentation pipeline as in SogCLR [Yuan et al., 2022], utilizing the torchvision implementation. This includes RandomResizedCrop (resizing to $224\times 224$), random ColorJitter, RandomGrayscale, random GaussianBlur, RandomHorizontalFlip, and normalization using ImageNet statistics.
@@
The network is pretrained for 200 epochs with a batch size of 128. We use the LARS optimizer [You et al., 2017] with a weight decay of $61e-$ 6 and momentum of 0.9. The temperature ($\tau$) is set to 0.1, and for SogCLR, $\gamma=0.9$. For GloFND, we set $\alpha=0.01$ and tune the starting epoch from $\{70,90,110\}$, after which we begin updating $\lambda_{i}$ with Adam updates using a learning rate of 0.05, $\beta_{1}=0.9$, and $\beta_{2}=0.98$. For FNC, $\alpha=0.01$ and the starting epoch is tuned from $\{10,30,50,70,90,110,130\}$, selecting the value that yields the best semi-supervised average performance.
Linear Evaluation.
@@
We evaluate GloFND’s ability to produce better representations through linear evaluation. Specifically, we freeze the encoder’s weights at the last iteration of pretraining, remove its projection head, and train a linear classifier (a single fully connected layer) on top of the encoder’s output.
@@
We train for 90, 285, 900, and 900 epochs corresponding to 100%, 10%, 1%, and 0.1% labeled data, respectively, with a batch size of 1024 and early stopping if the validation accuracy does not improve for 100 epochs. We use AdamW [Loshchilov and Hutter, 2019] with a weight decay of 0, momentum of 0.9, and a learning rate of 0.1. The same augmentation pipeline used in SogCLR is applied for linear evaluation. For training, we use RandomResizedCrop (resizing to $224\times 224$), RandomHorizontalFlip, and normalization. For testing, we resize the images to $256\times 256$, apply CenterCrop to $224\times 224$, and normalize.
Transfer Learning Datasets.
We additionally examine the transfer learning performance on Food-101 [Bossard et al., 2014], CIFAR-10 and CIFAR-100 [Krizhevsky, 2009], Stanford Cars [Krause et al., 2013], Describable Textures Dataset (DTD) [Cimpoi et al., 2014], Oxford-IIIT Pets [Parkhi et al., 2012], Caltech-101 [Li et al., 2022], and Oxford 102 Flowers [Nilsback and Zisserman, 2008]. We follow the evaluation protocols in the papers introducing these datasets, i.e., we report top-1 accuracy for Food-101, CIFAR-10, CIFAR-100, Stanford Cars, and DTD; and mean per-class accuracy for Oxford-IIIT Pets, Caltech-101, and Oxford 102 Flowers. We report results on the test set and, for DTD, we report results only for the first split. Caltech-101 defines no train/test split, so we randomly select 20% of images per class to create the test set.
Transfer Learning Evaluation.
We train an $\ell_{2}$-regularized multinomial logistic regression classifier on features extracted from the frozen pretrained network after removing the projector head. For each method, we select the pretrained network that achieved the highest semi-supervised average performance, as used in the semi-supervised results.
@@
Appendix H Additional Details for Bimodal Experiments
Datasets.
For bimodal learning, we use the Conceptual Captions 3M (CC3M) [Sharma et al., 2018] dataset. Because some links have expired, our downloaded training set of CC3M contains $2,723,840$ image-text pairs. We evaluate the performance by leveraging the Datacomp Benchmark [Gadre et al., 2023], which includes 38 zero-shot downstream tasks. We report the average performance, named Datacomp. For each scenario, we select the model with the best Datacomp average and also report its average performance on two subsets of the tasks: zero-shot image classification on ImageNet and its different variants (IN & Variants), and zero-shot cross-modal image-text retrieval. IN & Variants includes ImageNet-1k [Russakovsky et al., 2015] and 6 ImageNet distribution shift datasets (i.e., ImageNet-Sketch [Wang et al., 2019], ImageNet-V2 [Recht et al., 2019], ImageNet-A [Hendrycks et al., 2021b], ImageNet-O [Hendrycks et al., 2021b], ImageNet-R [Hendrycks et al., 2021a], and ObjectNet [Barbu et al., 2019]). Retrieval tasks consist of Flickr30K [Plummer et al., 2017], MSCOCO [Lin et al., 2015], and WinoGAViL [Bitton et al., 2022].
Experiment Setup.
Following previous work [Wei et al., 2024], we use a 12-layer transformer [Vaswani et al., 2017] as the text encoder and ResNet50 as the vision encoder. All experiments are conducted in a multi-node setting with 2 nodes, each equipped with two A100 40GB GPUs. We pretrain for 37 epochs with a global batch size of 1024. The AdamW [Loshchilov and Hutter, 2019] optimizer is used with $(\beta_{1},\beta_{2})=(0.9,0.999)$, $\epsilon=1e-8$, and a learning rate of $31e-$ 3.A weight decay of 0.1 is applied, with a warm-up period of 10k steps. The learning rate follows a cosine schedule, initially increasing linearly during the warm-up phase and then decreasing according to a cosine function. A cosine $\gamma$ schedule is employed, with a minimum $\gamma$ of 0.2 and decay epochs set to 18.

### v2

@@ 4.2 Bimodal Experiments
Experiment Setup. Following previous work (Wei et al., 2024), we use a 12-layer transformer (Vaswani et al., 2017) as the text encoder, and ResNet50 as the vision encoder. All experiments are conducted in a multi-node setting with 2 nodes, each with two A100 40GB GPUs. We pretrain for 37 epochs with a global batch size of 1024. For SogCLR, we start using FNC/GloFND after 15 epochs, setting $\alpha=5e-4$, while for FastCLIP we start after 20 epochs with $\alpha=1e-3$. For both losses, we initialize $\lambda_{i}=1$ and use Adam updates with a learning rate of 0.05. More details on hyperparameters can be found in Appendix D.2.
@@ 4.3 Ablation Study
Impact of Starting Epoch. GloFND requires a sufficiently pre-trained network to ensure that the similarity between embeddings reflects semantic similarity. This is achieved by applying GloFND after a certain number of training epochs. However, the optimal start time involves a trade-off. If GloFND is applied too early, the embedding space may not be well-formed, leading to incorrect classification of pairs as false negatives. Conversely, if applied too late, the potential benefits of GloFND may be limited due to insufficient training time. To assess the necessity and impact of this “wait” period, we evaluate GloFND’s semi-supervised performance on ImageNet100 as we vary the start epoch in $\{10,30,50,70,90,110\}$, keeping the total number of training epochs fixed at 200. The linear evaluation results, presented in LABEL:fig:warmup_100, show that GloFND’s performance improves as the wait time increases, peaking at 70 epochs. Beyond this point, performance begins to decline. FNC shows a similar trend, with its performance degrading when applied after 110 epochs.
Computational Efficiency. When using GloFND with a contrastive method based on embedding similarity (e.g., SimCLR, SogCLR, and CLIP), pairwise similarities are already computed as part of the loss function. Thus, the only additional computations required are: (1) updating the $\lambda_{i}$ values for the mini-batch samples, which involves a simple gradient computation (Equation 3.2.1), and (2) filtering the false negatives, which can be done through simple matrix operations. Both operations involve basic matrix computations and run in linear time with respect to the number of pairs in a batch ($O(B^{2})$, where $B$ is the batch size). The computational overhead of GloFND is minimal compared to the cost of cosine similarity computation and forward/backpropagation. Our experiments on ImageNet100 with a batch size of 128 show that GloFND introduces only a 2% increase in per-epoch training time for SogCLR (435.19 s for SogCLR + GloFND vs. 426.67 s for SogCLR). This overhead is comparable to that of the batch-wise method FNC (434.06 s).
@@ 5 Conclusions
Limitations. Since the focus of this paper is on false negative detection for contrastive learning, we address false negatives through filtering. While this straightforward approach has proven effective in our settings, future work could explore more advanced methods that may further enhance downstream performance. Additionally, the benefits of GloFND and similar false-negative techniques on downstream tasks depend on the proportion of false negatives in the pretraining dataset and how false negatives are defined within the downstream task.
Acknowledgments
VB, BW and TY were partially supported by National Science Foundation Award #2306572 and #2147253, National Institutes of Health Award #R01HL168116. CL was partially supported by National Institutes of Health Award #R01HL168116.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Proof of Lemma 3.1
Proof.
For simplicity, we denote that $n_{i}^{-}=|S_{i}^{-}|$ and $s_{j}$ is the $j$-th largest value in $\{\mathrm{sim}(\mathbf{z}_{i},E_{\mathbf{w}}(\mathbf{x}))\mid\mathbf{x}\in\mathcal{S}_{i}^{-}\}$, i.e., $s_{1}\geq s_{2}\geq\dotsc\geq s_{n_{i}^{-}}$. A subgradient $\phi_{i}^{\prime}(\lambda_{i})$ of the objective at $\lambda_{i}$ in (2) is
$\phi_{i}^{\prime}(\lambda_{i})=\alpha n_{i}^{-}-\sum_{j=1}^{n_{i}^{-}}\psi(s_{j}-\lambda_{i}),\quad\psi(s_{j}-\lambda_{i})=\begin{cases}1,&s_{j}>\lambda_{i}\\ \epsilon,&s_{j}=\lambda_{i}\\ 0,&s_{j}<\lambda_{i},\end{cases}$
where $\epsilon\in[0,1]$. We define $k=\lceil\alpha n_{i}^{-}\rceil$. $0\in\partial\phi_{i}(\lambda_{i})$ only happens when $\lambda_{i}\in[s_{k^{\prime\prime}},s_{k^{\prime}})$ since $k-1<\alpha n_{i}^{-}\leq k$, where $k^{\prime}=\max\{j\mid s_{j}>s_{k},j>k\}$, $k^{\prime\prime}=\min\{k^{\prime\prime\prime},k+1\}$, $k^{\prime\prime\prime}=\max\{j\mid s_{j}<s_{k},j>k\}$. Thus, $s_{k}$ is a solution to (2). Besides, when $\alpha n_{i}^{-}$ is an integer (i.e. $k=\alpha n_{i}^{-}=\lceil\alpha n_{i}^{-}\rceil$), any value between $[s_{k},s_{k+1}]$ is also a solution to (2), which could be different from $s_{k}$.
Appendix B More Details on Extension to Bimodal CL
@@
Our approach GloFND can be extended to resolve the global false negative discovery in bimodal CL, e.g., CLIP (Radford et al., 2021). Consider a dataset of image-text pairs $\mathcal{D}=\{(\mathbf{x}_{1},\mathbf{t}_{1}),\dotsc,(\mathbf{x}_{n},\mathbf{t}_{n})\}$, a collection of image augmentation operators $\mathcal{P}_{I}$, and a collection of text augmentation operators $\mathcal{P}_{T}$. Suppose that the encoder network $E_{I}$ for images and the encoder network $E_{T}$ for text are parametrized by $\mathbf{w}$. For each $(\mathbf{x}_{i},\mathbf{t}_{i})\in\mathcal{D}$, the negative dataset for each anchor image $\mathbf{x}_{i}$ is $\mathcal{S}_{I,i}^{-}=\{A_{T}(\mathbf{t})\mid\forall A_{T}\in\mathcal{P}_{T},\forall(\mathbf{x},\mathbf{t})\in\mathcal{D}\}$ while the negative dataset for each anchor text $\mathbf{t}_{i}$ is $\mathcal{S}_{T,i}^{-}=\{A_{I}(\mathbf{x})\mid\forall A_{I}\in\mathcal{P}_{I},\forall(\mathbf{x},\mathbf{t})\in\mathcal{D}\}$. For the $i$-th anchor image $\mathbf{x}_{i}$ with representation $\mathbf{z}_{I,i}$, the threshold $\lambda_{I,i}\in[-1,1]$ for finding the top $\alpha$text neighbors among all negatives $\mathcal{S}_{I,i}^{-}$ can be solved by
$\min_{\nu\in[-1,1]}\phi_{I,i}(\nu),$
$\phi_{I,i}(\nu)=\nu\alpha+\frac{1}{|\mathcal{S}_{I,i}^{-}|}\sum_{\mathbf{t}\in\mathcal{S}_{I,i}^{-}}\left(\mathrm{sim}(\mathbf{z}_{I,i},E_{T}(\mathbf{t}))-\nu\right)_{+}.$
@@
Similarly, we can obtain the threshold $\lambda_{T,i}\in[-1,1]$ for the $i$-th anchor text $\mathbf{t}_{i}$. Given the thresholds $\bm{\lambda}_{I},\bm{\lambda}_{T}$, the bimodal contrastive loss can be written as
@@
Appendix C High-level Intuition for $\alpha$ Hyperparameter
@@
The hyperparameter $\alpha$ allows GloFND to adapt to different definitions of false negatives, which are inherently dependent on the desired level of granularity. For example, in a dataset like ImageNet, consider two classification settings: (1) a coarse-grained task of classifying between cars and animals, and (2) fine-grained task of classifying dog breeds. Two images of a dog from different breeds might be considered a false negative in the coarse-grained case (1), but not in the fine-grained one (2). Consequently, the optimal percentage of false negatives, controlled by $\alpha$, varies with the chosen granularity. By adjusting $\alpha$, GloFND can flexibly align with different levels of semantic resolution, i.e., granularity.
@@
The value of $\alpha$ can be set based on prior knowledge (e.g., expected rate of false negatives or desired granularity of learned representations) or tuned like other hyperparameters such as the temperature. If $\alpha$ is too low, GloFND may fail to identify sufficient false negatives, leading to minimal impact on the learned representations, though not degrading performance, as setting $\alpha=0$ is equivalent to disabling GloFND. Conversely, if $\alpha$ is too high, GloFND may identify too many false negatives. If these are filtered out during training, the reduced number of negative pairs can limit contrastive learning, potentially harming performance. For tuning, it is recommended to start with a low $\alpha$ and gradually increase it until no further performance gains are observed.
Appendix D More Details on Experiments
@@
All the experiments are implemented using the PyTorch (Paszke et al., 2019) library. The unimodal experiments are run on a single NVIDIA A30 with 24GB memory size, while the bimodal experiments make use of a multi-node setup with 2 nodes, each with 2 NVIDIA A100 GPUs with 40GB each. The estimated amount of time to run a single experiment is 1.5 days for ImageNet100, and 14 hours for CC3M.
D.1 Additional Details for Unimodal Experiments
@@
Experiment Setup. Following prior work (Yuan et al., 2022), we pretrain a ResNet-50 (He et al., 2015) with a 2-layer $128\times 128$ projection head on top of the backbone encoder. We use square root learning rate scaling ($0.075\times\sqrt{\text{BatchSize}}$) with a cosine decay schedule without restart. Additionally, we apply a linear learning rate warm-up for 10 epochs, where the learning rate linearly increases to its maximum value.
@@
We adopt the same augmentation pipeline as in SogCLR (Yuan et al., 2022), utilizing the torchvision implementation. This includes RandomResizedCrop (resizing to $224\times 224$), random ColorJitter, RandomGrayscale, random GaussianBlur, RandomHorizontalFlip, and normalization using ImageNet statistics.
@@
The network is pretrained for 200 epochs with a batch size of 128. We use the LARS optimizer (You et al., 2017) with a weight decay of $61e-$ 6 and momentum of 0.9. The temperature ($\tau$) is set to 0.1, and for SogCLR, $\gamma=0.9$. For GloFND, we set $\alpha=0.01$ and tune the starting epoch from $\{70,90,110\}$, after which we begin updating $\lambda_{i}$ with Adam updates using a learning rate of 0.05, $\beta_{1}=0.9$, and $\beta_{2}=0.98$. For FNC, $\alpha=0.01$ and the starting epoch is tuned from $\{10,30,50,70,90,110,130\}$, selecting the value that yields the best semi-supervised average performance.
@@
Linear Evaluation. We evaluate GloFND’s ability to produce better representations through linear evaluation. Specifically, we freeze the encoder’s weights at the last iteration of pretraining, remove its projection head, and train a linear classifier (a single fully connected layer) on top of the encoder’s output.
@@
We train for 90, 285, 900, and 900 epochs corresponding to 100%, 10%, 1%, and 0.1% labeled data, respectively, with a batch size of 1024 and early stopping if the validation accuracy does not improve for 100 epochs. We use AdamW (Loshchilov & Hutter, 2019) with a weight decay of 0, momentum of 0.9, and a learning rate of 0.1. The same augmentation pipeline used in SogCLR is applied for linear evaluation. For training, we use RandomResizedCrop (resizing to $224\times 224$), RandomHorizontalFlip, and normalization. For testing, we resize the images to $256\times 256$, apply CenterCrop to $224\times 224$, and normalize.
Transfer Learning Datasets. We additionally examine the transfer learning performance on Food-101 (Bossard et al., 2014), CIFAR-10 and CIFAR-100 (Krizhevsky, 2009), Stanford Cars (Krause et al., 2013), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Oxford-IIIT Pets (Parkhi et al., 2012), Caltech-101 (Li et al., 2022), and Oxford 102 Flowers (Nilsback & Zisserman, 2008). We follow the evaluation protocols in the papers introducing these datasets, i.e., we report top-1 accuracy for Food-101, CIFAR-10, CIFAR-100, Stanford Cars, and DTD; and mean per-class accuracy for Oxford-IIIT Pets, Caltech-101, and Oxford 102 Flowers. We report results on the test set and, for DTD, we report results only for the first split. Caltech-101 defines no train/test split, so we randomly select 20% of images per class to create the test set.
Transfer Learning Evaluation. We train an $\ell_{2}$-regularized multinomial logistic regression classifier on features extracted from the frozen pretrained network after removing the projector head. For each method, we select the pretrained network that achieved the highest semi-supervised average performance, as used in the semi-supervised results.
@@
D.2 Additional Details for Bimodal Experiments
Datasets. For bimodal learning, we use the Conceptual Captions 3M (CC3M) (Sharma et al., 2018) dataset. Because some links have expired, our downloaded training set of CC3M contains $2,723,840$ image-text pairs. We evaluate the performance by leveraging the Datacomp Benchmark (Gadre et al., 2023), which includes 38 zero-shot downstream tasks. We report the average performance, named Datacomp. For each scenario, we select the model with the best Datacomp average and also report its average performance on two subsets of the tasks: zero-shot image classification on ImageNet and its different variants (IN & Variants), and zero-shot cross-modal image-text retrieval. IN & Variants includes ImageNet-1k (Russakovsky et al., 2015) and 6 ImageNet distribution shift datasets (i.e., ImageNet-Sketch (Wang et al., 2019), ImageNet-V2 (Recht et al., 2019), ImageNet-A (Hendrycks et al., 2021b), ImageNet-O (Hendrycks et al., 2021b), ImageNet-R (Hendrycks et al., 2021a), and ObjectNet (Barbu et al., 2019)). Retrieval tasks consist of Flickr30K (Plummer et al., 2017), MSCOCO (Lin et al., 2015), and WinoGAViL (Bitton et al., 2022).
Experiment Setup. Following previous work (Wei et al., 2024), we use a 12-layer transformer (Vaswani et al., 2017) as the text encoder and ResNet50 as the vision encoder. All experiments are conducted in a multi-node setting with 2 nodes, each equipped with two A100 40GB GPUs. We pretrain for 37 epochs with a global batch size of 1024. The AdamW (Loshchilov & Hutter, 2019) optimizer is used with $(\beta_{1},\beta_{2})=(0.9,0.999)$, $\epsilon=1e-8$, and a learning rate of $31e-$ 3.A weight decay of 0.1 is applied, with a warm-up period of 10k steps. The learning rate follows a cosine schedule, initially increasing linearly during the warm-up phase and then decreasing according to a cosine function. A cosine $\gamma$ schedule is employed, with a minimum $\gamma$ of 0.2 and decay epochs set to 18.